# On eigenfunctions of the block graphs of geometric Steiner systems

### v2

@@ Abstract
journal: Journal of Combinatorial Designs
@@ 1 Introduction
This paper lies in the context of the studies of eigenfunctions of graphs having minimum cardinality of support (see SV21 for the background). One of the tools is the weight-distribution bound, a lower bound on the cardinality of support of an eigenfunction of a distance-regular graph corresponding to a non-principal eigenvalue. Note that the tightness of the weight-distribution bound was previously shown in general for the smallest eigenvalue of a Grassmann graph $J_{q}(n+1,e)$ (see (KMP16, , Section 4)). However, a characterisation of optimal eigenfunctions was not obtained. Motivated by this open problem, we consider the class of strongly regular Grassmann graphs and give the required characterisation in this case.
@@
Similar to (BR98, , Theorem 2.4.4), the following proposition gives an explicit general algebraic description of an affine regulus in $\operatorname{AG}(n,q)$.
@@
In this paper we are not only interested in optimal eigenfunctions. We also discuss an approach for studying equitable 2-partitions based on $(1,-1,0)$-valued eigenfunctions. It is well-known that the investigation of equitable $2$-partitions is equivalent to the investigation of eigenfunctions with two values (see Lemma 7, for example). Since eigenfunctions corresponding to distinct eigenvalues of a graph are orthogonal, one can use $\theta_{1}$-eigenfunctions of a regular graph $G$ to restrict possible $\theta_{2}$-equitable $2$-partitions of the graph $G$ whenever $\theta_{1}\neq\theta_{2}$ (see T20 for a general discussion).
@@
Let $\theta$ be the positive non-principal eigenvalue of a Grassmann graph $J_{q}(4,2)$. A part of a $\theta$-equitable 2-partitions of $J_{q}(4,2)$ is called a Cameron-Liebler line class. It is well-known that one of the equivalent definitions of a Cameron-Liebler line class is as follows (see (P91, , Theorem 1(v))): a class of lines $L$ in $\operatorname{PG}(3,q)$ is Cameron-Liebler if and only if $|L\cap R|=|L\cap R^{opp}|$ for all reguli $R$ in $\operatorname{PG}(3,q)$. We note that this condition naturally follows from Theorem 1 and Corollary 1. Cameron-Liebler lines classes in affine spaces were recently studied in DMSS20 and DIMS21 .
@@ 2.1 The weight-distribution bound for strongly regular graphs
holds, where $G(u)$ is the set of neighbours of the vertex $u$. Although eigenfunctions of graphs receive less attention of researchers in contrast to their eigenvalues, there are still plenty of related literature. We refer to the recent survey SV21 for a summary of results on the problem of finding the minimum cardinality of support of eigenfunctions of graphs and characterising the optimal eigenfunctions. For an eigenspace of a graph, an eigenfunction is called optimal if it has the minimum cardinality of support among all eigenfunctions in this eigenspace.
The following lemma gives a lower bound for the number of non-zeroes (i.e., the cardinality of the support) for an eigenfunction of a strongly regular graph. In fact, this is a special case of a more general result for distance-regular graphs (KMP16, , Section 2.4).
@@
If follows from (KMP16, , Corollary 1) and the fact that a primitive strongly regular graph with parameters $(v,k,\lambda,\mu)$ is a distance-regular graph of diameter $2$ with $a_{1}=\lambda$, $b_{0}=k$ and $c_{2}=\mu$. ∎
@@
Suppose $G$ is strongly regular with parameters $(v,k,\lambda,\mu)$. Note that $s,r$ are the two solution of the quadratic equation $x^{2}-(\lambda-\mu)x-(k-\mu)=0$ according to Lemma 1. Thus, $(s-\lambda)s-k=-\mu(s+1)$ and $(r-\lambda)r-k=-\mu(r+1)$. It follows that
@@
A clique in a strongly regular graph $G$ is called Delsarte clique if its size meets the Delsarte-Hoffman bound (see (BCN89, , Proposition 1.3.2)).
@@
It follows from (KMP16, , Theorem 3, Theorem 4) and the fact that taking the complement of a strongly regular graph preserves the eigenspaces corresponding to the non-principal eigenvalues, which describes the support of an eigenfunction of a strongly regular graph that meets the weight-distribution bound. ∎
@@ 2.2 Geometric Steiner systems
Let $V(n,q)$ be a vector space of dimension $n$ over a finite field $\mathbb{F}_{q}$. The affine space $\operatorname{AG}(n,q)$ is the geometry whose points, lines, planes, etc., are the cosets of the subspaces of $V(n,q)$ of dimension $0,1,2,\ldots,n$. Denote by $AS(n,q)$ the Steiner system whose points and blocks are the points and the lines of the affine space $\operatorname{AG}(n,q)$, and let us call such a Steiner system an affine Steiner system. Such a system is known to be a 2-$(q^{n},q,1)$ design and is a called an affine design throughout the paper.
The projective space $\operatorname{PG}(n,q)$ is the geometry whose points, lines, planes, etc., are the subspaces of $V(n+1,q)$ of dimension $11,2,\ldots,n+$ 1. Denote by $PS(n,q)$ the Steiner system whose points and blocks are the points and the lines of the projective space $\operatorname{PG}(n,q)$, and let us call such a Steiner system a projective Steiner system. Such a system is known to be a 2-$((q^{n}-1)/(q-1),q+1,1)$ design and is a called a projective design throughout the paper.
@@
The following proposition gives the weight-distribution bound for the smallest eigenvalue of the block graph of a 2-$(N,M,1)$ designs.
@@ 2.3 Reguli in projective spaces
In this section, we follow the definitions from BR98 . Let $P$ be a projective space. A set $S$ of subspaces of $P$ is called skew if no two distinct subspaces of $S$ have a point in common. Let $S$ be a set of skew subspaces. A line is called a transversal of $S$ if it intersects each subspace of $S$ in exactly one point.
Lemma 4 ((BR98, , Lemma 2.4.1)).
@@
It follows from the definition that the set $R^{opp}$ of all transversals of a regulus $R$ again form a regulus; we call it the opposite regulus of $R$.
@@ 2.4 Equitable partitions as eigenfunctions
Let $G$ be a $k$-regular graph with the vertex set $V(G)$. Let $\Pi:=(V_{1},\ldots,V_{t})$ be a partition of $V(G)$ into $t$ parts ($t$-partition). The partition $\Pi$ is said to be equitable $t$-partition if for any $i,j\in\{1,\ldots,t\}$ there is a constant $p_{ij}$ such that any vertex from the part $V_{i}$ is adjacent to precisely $p_{ij}$ vertices from the part $V_{j}$. The square matrix $P_{\Pi}:=(p_{ij})_{i,j=1}^{t}$ is called the quotient matrix of the equitable $t$-partition $\Pi$. Since all row sums of the adjacency matrix $A$ of $G$ and the quotient matrix $P_{\Pi}$ are equal to $k$, both matrices have eigenvalue $k$. Moreover, it follows from (GR01, , Theorem 9.3.3), that every eigenvalue of $P$ is an eigenvalue of $A$. The eigenvalue $k$ of $A$ is called principal. An eigenvalue $\theta$ of $A$ is called non-principal if $\theta\neq k$. If $\Pi$ is an equitable $2$-partition, then precisely one non-principal eigenvalue $\theta$ is an eigenvalue of the quotient matrix $P_{\Pi}$. In this case we say that the equitable $2$-partition $\Pi$ is $\theta$-equitable.
@@ 3.1 Proof of Theorem 1
In this section we consider the block graphs of projective Steiner system $PS(n,q)$, known as the class of strongly regular Grassmann graphs $J_{q}(n+1,2)$. A Grassmann graph $J_{q}(n+1,2)$ is known to be arc-transitive and have Delsarte cliques (the pencil of lines through a point is an example of a Delsarte clique). In view of Lemma 3(3), any copy (as an induced subgraph) of the complete bipartite graph with parts of size $q+1$ in $J_{q}(n+1,2)$ gives rise to an $-(q+1)$-eigenfunction of $J_{q}(n+1,2)$ whose cardinality of support meets the weight-distribution bound. It follows from Lemma 5(2) that the reguli in all 3-dimensional subspaces of $\operatorname{PG}(n,q)$ exhaust the induced complete bipartite subgraphs with parts of size $q+1$ in $J_{q}(n+1,2)$, which gives the required characterisation of optimal $-(q+1)$-eigenfunctions. This proves Theorem 1.
@@ 3.2 Proof of Theorem 2, Proposition 1 and Proposition 2
In this section we consider the block graphs of affine Steiner system $AS(n,q)$, denoted by $X_{q}(n,1)$. These graphs are known to be arc-transitive and have Delsarte cliques (the pencil of lines through a point is an example of a Delsarte clique). In view of Lemma 3(3), any copy (as an induced subgraph) of the complete bipartite graph with parts of size $q$ in $X_{q}(n,1)$ gives rise to an $-q$-eigenfunction of $X_{q}(n,1)$ whose cardinality of support meets the weight-distribution bound.
@@
To complete the proof, let us show that if $\ell_{1},\ell_{2},\ell_{3}$ is a triple of lines from Case 2, then this triple cannot be extended to an affine regulus. Suppose to the contrary there exist lines $m_{1},\ldots,m_{q-3}$ such that $\ell_{1},\ell_{2},\ell_{3},m_{1},\ldots,m_{q-3}$ is an affine regulus. Consider the projective regulus $\ell_{1}^{\prime},\ell_{2}^{\prime},\ell_{3}^{\prime},k_{1}^{\prime},\ldots,k_{q-2}^{\prime}$ uniquely determined by the triple $\ell_{1}^{\prime},\ell_{2}^{\prime},\ell_{3}^{\prime}$. Note that each transversal to this projective regulus does not lie in the plane at infinity (otherwise, the triple $\ell_{1},\ell_{2},\ell_{3}$ would be from Case 1), that is, each such a transversal intersects the plane at infinity at one point. Note that the restriction of each such a transversal to the affine part is a transvesral to the affine regulus. Thus, the affine regulus has at least $q+1$ transversals. By the pigeonhole principle, there exists a point of the affine regulus incident to two transversals, which means the entire affine regulus lies in the affine plane determined by these two transversals, a contradiction.

### v3

@@ Abstract
journal: Arxiv
@@ 1 Introduction
This paper lies in the context of the studies of eigenfunctions of graphs having minimum cardinality of support (see [11] for the background). One of the tools is the weight-distribution bound, a lower bound on the cardinality of support of an eigenfunction of a distance-regular graph corresponding to a non-principal eigenvalue. Note that the tightness of the weight-distribution bound was previously shown in general for the smallest eigenvalue of a Grassmann graph $J_{q}(n+1,e)$ (see [9, Section 4]). However, a characterisation of optimal eigenfunctions was not obtained. Motivated by this open problem, we consider the class of strongly regular Grassmann graphs and give the required characterisation in this case.
@@
Similar to [2, Theorem 2.4.4], the following proposition gives an explicit general algebraic description of an affine regulus in $\operatorname{AG}(n,q)$.
@@
In this paper we are not only interested in optimal eigenfunctions. We also discuss an approach for studying equitable 2-partitions based on $(1,-1,0)$-valued eigenfunctions. It is well-known that the investigation of equitable $2$-partitions is equivalent to the investigation of eigenfunctions with two values (see Lemma 7, for example). Since eigenfunctions corresponding to distinct eigenvalues of a graph are orthogonal, one can use $\theta_{1}$-eigenfunctions of a regular graph $G$ to restrict possible $\theta_{2}$-equitable $2$-partitions of the graph $G$ whenever $\theta_{1}\neq\theta_{2}$ (see [12] for a general discussion).
@@
Let $\theta$ be the positive non-principal eigenvalue of a Grassmann graph $J_{q}(4,2)$. A part of a $\theta$-equitable 2-partition of $J_{q}(4,2)$ is called a Cameron-Liebler line class. It is well-known that one of the equivalent definitions of a Cameron-Liebler line class is as follows (see [10, Theorem 1(v)]): a class of lines $L$ in $\operatorname{PG}(3,q)$ is Cameron-Liebler if and only if $|L\cap R|=|L\cap R^{opp}|$ for all reguli $R$ in $\operatorname{PG}(3,q)$. We note that this condition naturally follows from Theorem 1 and Corollary 1. Cameron-Liebler classes have been intensively studied (see, for example, [6] and [5]). Cameron-Liebler lines classes in affine spaces have been recently studied in [4] and [3].
@@ 2.1 The weight-distribution bound for strongly regular graphs
holds, where $G(u)$ is the set of neighbours of the vertex $u$. Although eigenfunctions of graphs receive less attention of researchers in contrast to their eigenvalues, there are still plenty of related literature. We refer to the recent survey [11] for a summary of results on the problem of finding the minimum cardinality of support of eigenfunctions of graphs and characterising the optimal eigenfunctions. For an eigenspace of a graph, an eigenfunction is called optimal if it has the minimum cardinality of support among all eigenfunctions in this eigenspace.
The following lemma gives a lower bound for the number of non-zeroes (i.e., the cardinality of the support) for an eigenfunction of a strongly regular graph. In fact, this is a special case of a more general result for distance-regular graphs [9, Section 2.4].
@@
If follows from [9, Corollary 1] and the fact that a primitive strongly regular graph with parameters $(v,k,\lambda,\mu)$ is a distance-regular graph of diameter $2$ with $a_{1}=\lambda$, $b_{0}=k$ and $c_{2}=\mu$. ∎
@@
Suppose $G$ is strongly regular with parameters $(v,k,\lambda,\mu)$. Note that $s,r$ are the two solutions of the quadratic equation $x^{2}-(\lambda-\mu)x-(k-\mu)=0$ according to Lemma 1. Thus, $(s-\lambda)s-k=-\mu(s+1)$ and $(r-\lambda)r-k=-\mu(r+1)$. It follows that
@@
A clique in a strongly regular graph $G$ is called a Delsarte clique if its size meets the Delsarte-Hoffman bound (see [1, Proposition 1.3.2]).
@@
It follows from [9, Theorem 3, Theorem 4] and the fact that taking the complement of a strongly regular graph preserves the eigenspaces corresponding to the non-principal eigenvalues, which describes the support of an eigenfunction of a strongly regular graph that meets the weight-distribution bound. ∎
@@ 2.2 Geometric Steiner systems
Let $V(n,q)$ be a vector space of dimension $n$ over a finite field $\mathbb{F}_{q}$. The affine space $\operatorname{AG}(n,q)$ is the geometry whose points, lines, planes, etc., are the cosets of the subspaces of $V(n,q)$ of dimension $0,1,2,\ldots,n$. Denote by $AS(n,q)$ the Steiner system whose points and blocks are the points and the lines of the affine space $\operatorname{AG}(n,q)$, and let us call such a Steiner system an affine Steiner system. Such a system is known to be a 2-$(q^{n},q,1)$ design and is called an affine design throughout the paper.
The projective space $\operatorname{PG}(n,q)$ is the geometry whose points, lines, planes, etc., are the subspaces of $V(n+1,q)$ of dimension $11,2,\ldots,n+$ 1. Denote by $PS(n,q)$ the Steiner system whose points and blocks are the points and the lines of the projective space $\operatorname{PG}(n,q)$, and let us call such a Steiner system a projective Steiner system. Such a system is known to be a 2-$((q^{n+1}-1)/(q-1),q+1,1)$ design and is a called a projective design throughout the paper.
@@
The following proposition gives the weight-distribution bound for the smallest eigenvalue of the block graph of a 2-$(N,M,1)$ design.
@@ 2.3 Reguli in projective spaces
In this section, we follow the definitions from [2]. Let $P$ be a projective space. A set $S$ of subspaces of $P$ is called skew if no two distinct subspaces of $S$ have a point in common. Let $S$ be a set of skew subspaces. A line is called a transversal of $S$ if it intersects each subspace of $S$ in exactly one point.
Lemma 4 ([2, Lemma 2.4.1]).
@@
It follows from the definition that the set $R^{opp}$ of all transversals of a regulus $R$ again forms a regulus; we call it the opposite regulus of $R$.
@@ 2.4 Equitable partitions as eigenfunctions
Let $G$ be a $k$-regular graph with the vertex set $V(G)$. Let $\Pi:=(V_{1},\ldots,V_{t})$ be a partition of $V(G)$ into $t$ parts ($t$-partition). The partition $\Pi$ is said to be anequitable $t$-partition if for any $i,j\in\{1,\ldots,t\}$ there is a constant $p_{ij}$ such that any vertex from the part $V_{i}$ is adjacent to precisely $p_{ij}$ vertices from the part $V_{j}$. The square matrix $P_{\Pi}:=(p_{ij})_{i,j=1}^{t}$ is called the quotient matrix of the equitable $t$-partition $\Pi$. Since all row sums of the adjacency matrix $A$ of $G$ and the quotient matrix $P_{\Pi}$ are equal to $k$, both matrices have eigenvalue $k$. Moreover, it follows from [8, Theorem 9.3.3], that every eigenvalue of $P$ is an eigenvalue of $A$. The eigenvalue $k$ of $A$ is called principal. An eigenvalue $\theta$ of $A$ is called non-principal if $\theta\neq k$. If $\Pi$ is an equitable $2$-partition, then precisely one non-principal eigenvalue $\theta$ is an eigenvalue of the quotient matrix $P_{\Pi}$. In this case we say that the equitable $2$-partition $\Pi$ is $\theta$-equitable.
@@ 3.1 Proof of Theorem 1
In this section we consider the block graphs of the projective Steiner system $PS(n,q)$, known as the class of strongly regular Grassmann graphs $J_{q}(n+1,2)$. A Grassmann graph $J_{q}(n+1,2)$ is known to be arc-transitive and have Delsarte cliques (the pencil of lines through a point is an example of a Delsarte clique). In view of Lemma 3(3), any copy (as an induced subgraph) of the complete bipartite graph with parts of size $q+1$ in $J_{q}(n+1,2)$ gives rise to an $-(q+1)$-eigenfunction of $J_{q}(n+1,2)$ whose cardinality of support meets the weight-distribution bound. It follows from Lemma 5(2) that the reguli in all 3-dimensional subspaces of $\operatorname{PG}(n,q)$ exhaust the induced complete bipartite subgraphs with parts of size $q+1$ in $J_{q}(n+1,2)$, which gives the required characterisation of optimal $-(q+1)$-eigenfunctions. This proves Theorem 1.
@@ 3.2 Proof of Theorem 2, Proposition 1 and Proposition 2
In this section we consider the block graphs of the affine Steiner system $AS(n,q)$, denoted by $X_{q}(n,1)$. These graphs are known to be arc-transitive and have Delsarte cliques (the pencil of lines through a point is an example of a Delsarte clique). In view of Lemma 3(3), any copy (as an induced subgraph) of the complete bipartite graph with parts of size $q$ in $X_{q}(n,1)$ gives rise to an $-q$-eigenfunction of $X_{q}(n,1)$ whose cardinality of support meets the weight-distribution bound.
@@
To complete the proof, let us show that if $\ell_{1},\ell_{2},\ell_{3}$ is a triple of lines from Case 2, then this triple cannot be extended to an affine regulus. Suppose to the contrary there exist lines $m_{1},\ldots,m_{q-3}$ such that $\ell_{1},\ell_{2},\ell_{3},m_{1},\ldots,m_{q-3}$ is an affine regulus. Consider the projective regulus $\ell_{1}^{\prime},\ell_{2}^{\prime},\ell_{3}^{\prime},k_{1}^{\prime},\ldots,k_{q-2}^{\prime}$ uniquely determined by the triple $\ell_{1}^{\prime},\ell_{2}^{\prime},\ell_{3}^{\prime}$. Note that each transversal to this projective regulus does not lie in the plane at infinity (otherwise, the triple $\ell_{1},\ell_{2},\ell_{3}$ would be from Case 1), that is, each such a transversal intersects the plane at infinity at one point. Note that the restriction of each such a transversal to the affine part is a transversal to the affine regulus. Thus, the affine regulus has at least $q+1$ transversals. By the pigeonhole principle, there exists a point of the affine regulus incident to two transversals, which means the entire affine regulus lies in the affine plane determined by these two transversals, a contradiction.